### v1

# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

Abstract.
We study compact hyperbolic surfaces and multiplicative observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil-Petersson random surfaces of large genus. This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves serve as large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and de Vèrdiere, thereby providing a more complete picture of the asymptotic behaviour of observables in the large-scale limit. Our approach does not rely on ball averaging operators or Nevo’s ergodic theorem. Instead, we introduce a new method based on the hyperbolic wave equation and the quantitative exponential mixing of the geodesic flow established by Ratner and Matheus.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University
1. Introduction
1.1. Background
In quantum systems, the state of the system can be described in terms of the eigenfunctions of its energy operator. Their evolution is governed by the Schrödinger equation, the quantum analogue of Newton’s equations of motion. Observables correspond to measurable quantities of the system. Understanding the interplay between the energy operator, its eigenfunctions, and the observables leads to a deeper understanding of the dynamics of the system. A central topic in quantum chaos is the study of localization and delocalization phenomena for the eigenfunctions of the energy operator. In this paper, the Hamiltonian system is assumed to be the geodesic flow $\phi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace-Beltrami operator on the manifold. For more background and a better overview, we refer to the following surveys: [Mar06, Zel05, Sar11, Zel19, Ana22].
Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator representing an observable. We are interested in the diagonal elements $\langle\psi_{j},A\psi_{j}\rangle$, which correspond to the expected value of $A$ in the state $\psi_{j}$, and in the off-diagonal elements $\langle\psi_{j},A\psi_{k}\rangle$, which represent transition amplitudes from $\psi_{j}$ to $\psi_{k}$ induced by the observation of $A$.
Quantum ergodicity theorem (Shnirelman [Shn74], Zelditch [Zel87], and Colin de Verdière [Col85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem. For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy
$\langle\psi_{j_{l}},A\psi_{j_{l}}\rangle\xrightarrow{l\to\infty}\omega(A),$
where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit cotangent bundle of $M$. Roughly speaking, the theorem asserts that the probability distributions induced by the eigenfunctions converge to the uniform distribution in the weak sense of testing the distributions against observables.
Shnirelman’s theorem concerns the diagonal elements of observables. Zelditch [Zel90, Zel96b] extended it to the off-diagonal elements, given that the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
$\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j},A\psi_{k}\right\rangle\right|^{2}<\epsilon.$
Informally, this means that the summed transition amplitudes between states separated approximately by energy $\tau$ decay to zero for almost all subsequences of eigenfunctions as the eigenvalue increases.
A unified formulation encompassing both of these results can be given following the works of multiple authors: [Shn74, Col85, Zel87, Laz93, Zel90, GL93, Zel96a, Zel96b, ZZ96, Sun97]. The version below adapts Theorems 4.1 and 4.4 from [Zel05].
Theorem 1.1 (Semi-classical quantum mixing, [Sun97, Zel05]).
Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenvalues $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator. Consider the following properties:
  • (i)
    $\lim_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\big|\langle\psi_{j},A\psi_{j}\rangle-\omega(A)\big|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{\lambda_{j}<\lambda}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j},A\psi_{k}\rangle|^{2}<\epsilon.$
Then the geodesic flow $\varphi_{t}$ is ergodic on $(S^{*}M,d\mu)$ if and only if properties (i) and (ii) hold. It is weakly mixing on $(S^{*}M,d\mu)$ if and only if properties (i), (ii), and (iii) hold. Here $S^{*}M$ denotes the unit tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
This theorem exhibits a delicate connection between the classical and quantum settings. If properties (i) and (ii) hold in the quantum setting, the corresponding classical geodesic flow is ergodic, and conversely. Likewise, if Properties (i)(iii) hold, the classical flow is weak mixing, and vice versa. We therefore say that a quantum system satisfying Properties (i) and (ii) is quantum ergodic, while one satisfying Properties (i)(iii) is quantum mixing.
Our focus will be on the compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature. In this context, the following theorem of Ratner [Rat87], with quantitative bounds due to Matheus [Mat13], will be central. The theorem is stated here in a slightly modified form using only the physical space observables.
Theorem 1.2 (Position space exponential mixing).
Let $X$ be a compact connected hyperbolic surface with geodesic flow $\phi_{t}$ and with $\lambda_{1}(X)$ being the smallest eigenvalue of the Laplace-Beltrami operator $\Delta$. Let $f,g\in L^{2}(X)$. Then for all $t\geq 0$,
$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(T^{1}X)}-\int_{T^{1}X}f\,dx\int_{T^{1}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4},\end{cases}$
and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $T^{1}X$, the unit tangent bundle.
The more general phase space version (see [Rat87, Mat13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing with exponential decay which further implies ergodicity. By Theorem 1.1, a compact connected hyperbolic surface satisfies properties (i)(iii), and hence is both quantum ergodic and quantum weak mixing.
Remark 1.3.
From Theorem 1.1, we see that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum weakly mixing. However, if we restrict the observables to be taken from the collection of the physical-space observables, i.e. the multiplicative operators corresponding to functions, then a $2$-torus with irrational aspect ratio is quantum ergodic (though still not quantum weak mixing). The $2$-sphere remains neither quantum ergodic nor quantum weak mixing. These examples can be analyzed explicitly due to the availability of closed-form expressions for their eigenfunctions and eigenvalues. An illustarting example showing this for a discrete circle with $N$ elements can be acquired using observable $a(x)=e^{\frac{2\pi i\,x}{N}}$; this was pointed out by Mostafa Sabri. This idea should translate to the $2$-dimensional surfaces as well. We see that the choice of observables matters.
These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other contexts, such as discrete settings. A pseudo-analogue of the above corollary to Theorem 1.1 was obtained for certain finite regular graphs by Anantharaman and Le Masson [AL15]. Their result, however, is not a strict analogue, since the high-energy limit is not meaningful for finite graphs, which possess only finitely many eigenvalues. Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$, as opposed to a fixed structure with increasing energy as in Shnirelman’s theorem. A special case of their result states that, under suitable assumptions, if $(a_{n})$ is a sequence of diagonal observables acting on functions on the vertices and if $(\Delta_{n})$ is a sequence of graph Laplacians each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then
$\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}\xrightarrow{n\to\infty}0.$
As in Shnirelman’s theorem, the averages of the expectation values converge to the uniform average. We refer to such results, in which the structure grows while energy remains bounded, as large-scale results, in contrast to the original results in which the energy frows. We refer to the latter results as large-energy results. The large-scale quantum ergodicity of graphs and related theorems is an active field with a vigorous interplay with the continuous settings. For more realted results in the discrete graph setting, see [DP12, BL13, EKYY13, TVW13, AL15, Gei15, BLL16, Ana17, AS17, BHKY17, BKY17, AS19, BS19, BHY19, McK22, MS23]. An especially impressive result establsihed in [BHKY17, BHY19, BKY17] shows a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] for random regular graphs; the randomization effectively suppresses pathological cases helping to establish stronger results compared to the deterministic setting.
Since the large-scale graph quantum ergodicity theorem is not a true analogue of Shnirelman’s theorem, it is natural to seek a corresponding surface result highlighting the distinctions from the original large-energy setting. Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for details). Their result resembles closely that of Anantharaman and Le Masson, with minor modifications reflecting the transition from discrete graphs to continuous surfaces.
Despite the distinct nature of the two limits, there exists a heuristic correspondence between the large-energy and large-scale settings. In the large-energy regime, as the energy grows, the eigenfunctions exhibit shorter wavelengths, resolving increasingly fine structures. In the large-scale regime, as the structure grows, finer details become accessible within a fixed energy window. Hence, these two asymptotic regimes may be regarded as analogous in spirit.
To our knowledge, the only existing large-scale off-diagonal results establishing analogues of properties (i), (ii), and (iii) of Theorem 1.1 (see Subsection 1.2) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices. Their setting resembles that of Anantharaman and Le Masson, but with graphs replaced by random matrices and observables by general matrices. Their result is remarkably strong as they require no averaging: they prove that
$\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|$
decays to zero at the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, thereby establishing a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94], and in fact extending it to the off-diagonal elements. This result is relevant as the random matrix theory is closely related to the quantum ergodicity on graphs and surfaces, especially if Berry’s conjecture (see [Ber77]) holds. For more related results for random matrices, see [ESY09a, ESY09b, BY17].
A new upcoming result by Bordenave, Letroit, and Sabri [BLS] establishes a quantum ergodicy and quantum mixing results on growing Schreier graphs where even the quantum ergodicy was previously unknown. To our knowledge, this is first mixing result in the graph setting. This is especially interesting as graphs are intuitively closer to the original setting on manifolds when compared to random matrices.
1.2. Main theorems
We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplicative operators, thereby complementing the result of Le Masson and Sahlsten [LS17]. We also prove a probabilistic analogue of this, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24]. The small difference is that we use the quantum Hamiltonian $\sqrt{-\Delta-1/4}$ instead of $\sqrt{-\Delta}$ but causes only a spectral shift. We already refer to Section 3 for more technical overview of the proof ideas as well as the novel ideas of the proofs used.
The proof follows the ideas of Le Masson and Sahlsten in that we establish a quantitative version which yields limiting forms of the theorem. We again divide the argument into a spectral and a geometric component. However, we employ a different propagator, namely the hyperbolic wave propagator
$P_{t}=h_{t}\!\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad h_{t}(x)=x^{-1}\sin(tx).$
The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=2^{-1/2}\,P_{t}\,u_{0}$ with initial data
$u(z,0)=0,\qquad u_{t}(z,0)=2\pi f,$
for some function $f$. For details, see [LP82]. A convenient form of the integral kernel of $P_{t}$ is established in Section 5.
Unlike the ball-averaging operator used by Le Masson and Sahlsten, which has no immediate connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner. This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten.
To access the off-diagonal elements $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we also introduce a modified propagator: $P_{t,\tau}=\cos(t\tau)\,P_{t}$. Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behaviour of the off-diagonal elements. This, however, complicates both the spectral and geometric analyses compared to the quantum ergodicity case, which concerns only the diagonal elements. In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, appearing in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$. The physical interpretation of this operator is not immediately transparent, but it can be verified that
$P_{t,\tau}^{*}aP_{t}=P_{t,\tau/2}^{*}aP_{t,\tau/2}+P_{-t,\tau/2}^{*}aP_{-t,\tau/2},$
meaning it equals the sum of two Heisenberg evolutions of $a$ with $P_{t,\tau/2}$, one forward and one backward in time.
Finally, we replace rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics. The idea of using the exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [HLM+]. The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result. Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum weak mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale. The more technical approach taken in the proofs is presented in Section 3.
We now state the main theorems.
Theorem 1.4 (Large-scale quantum ergodicity and quantum mixing).
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Assume that the sequence has the following properties:
  • (BSC)
    Benjamini–Schramm convergence: for every $R>0$,
    $\lim_{n\to\infty}\frac{\big|\{x\in X_{n}:\operatorname{InjRad}_{X_{n}}(x)<R\}\big|}{|X_{n}|}=0,$
  • (EXP)
    Expander property: the sequence admits a uniform lower bound on the spectral gap,
  • (UND)
    Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
  • (i)
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon.$
Properties (i)(iii) are the large-scale analogues of Properties (i)(iii) in Theorem 1.1. Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying (i)(iii) as large-scale quantum mixing. As in Theorem 1.1, the choice of observables plays a role, recall the example on the unit circle where quantum mixing fails discussed in the earlier remark, Remark 1.3. We note that this idea can be modified to prove that the $2$-torii of increasing area are not quantum mixing for physical-space observables, though there are sequences of $2$-torii that are, interestingly, quantum ergodic using the explicit formula of the eigenfunctions. We refer to the earlier illustrating example discussed This observation also illustrates that large-scale quantum ergodicity and large-scale quantum weak mixing are properties specific to certain sequences of hyperbolic surfaces rather than to general sequences of Riemannian manifolds.
Le Masson and Sahlsten [LS17] have already proven that property (i) holds in this setting. Hence to establish Theorem 1.4, it remains to prove properties (ii) and (iii). To prove the previous theorem, we need the following quantitative version of it.
Theorem 1.5 (Quantitative version).
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
$0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
and fix $\tau\in\mathbb{R}$. Then
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The coefficient appearing is
$C_{8}=50000\pi.$
Remark 1.6.
We observe from the statement of Theorem 1.5 that in the proof of Theorem 1.4 in Section 8, we could relax many of the conditions. E.g. we could let the spectral gap and the injectivity radius to shrink slowly as we proceed in the sequence. Thereby Theorem 1.4 could be extended to encompass a larger family of sequences.
An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces. This follows from Katz et al. [KSV07], who showed that the injectivity radius grows (ensuring properties (BSC) and (UND)), and from Kim et al. [KSR03], who established a uniform lower bound on the spectral gap (ensuring property (EXP)). The result of Kim et al. is the best deterministic bound currently available. The first, weaker, bound is Selberg’s $3/16$ bound, see [Sel65].
Moreover, the class of admissible sequences is quite broad which is seen when using probabilistic approach using the Weil-Petersson random model; for background, see [Mir13, Wri20]. By Le Masson and Sahlsten [LS24], for a sequence $(X_{g})$ of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is the Weil-Petersson probability measure on the moduli space $\mathcal{M}_{g}$ of hyperbolic surfaces of genus $g$, the condition (i) of Theorem 1.4 holds with high probability as $g$ grows to infinity. We state now a probabilistic analogue of Theorem 1.4, thus showing that a growing Weil-Petersson random surface sequence is quantum ergodic and quantum mixing with high probability and further complementing the results of Le Masson and Sahlsten [LS24].
Theorem 1.7 (Probabilistic version).
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
As mentioned, Property (i) was proved by Le Masson and Sahlsten [LS24]. Hence it remains to prove Properties (ii) and (iii) where the appraoch of the proof follows closely that of Theorem 1.4. However, as the assumptions are lighter in the probabilistic setting we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law, Theorem 8.1. We need analogues of these in the proof of Theorem 1.7. Theorem 8.2 consisting of recent results in the random surface theory offers us the required tools.
The use of Weil-Petersson random surfaces in quantum ergodicity is currently very active. Randomization effectively suppresses pathological cases helping to establish stong results that the analogous deterministic settings are far away from acquiring. For results in Benjamini-Schramm convergence and injectivity radius in Weyl-Peterson random model, see [Mir13, Mon21]. For recent results on the spectral gap, see [AM24, LW24, Mag24, HMT25, HMN25, WX25]. There are other random models and we especially refer to late groundbreaking results using the random covering model, see [MN20, MNP22, HM23]. See an interesting application of Property (EXP) to the Bose-Einstein condensation of random hyperbolic surfaces [LS22].
1.3. Further discussions
Like Le Masson and Sahlsten [LS17], we restrict the energy window to a compact interval in $\left(1/4,\infty\right)$. This assumption plays a crucial role in our proofs. However, it is plausible that with a different treatment of the low-lying modes—namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in [0,1/4)[0,1/4)—one could extend the results to an energy window covering the full range $0,\infty)$.
In contrast to Theorem 1.1, Theorem 1.4 is not a characterization-type result linking the classical and quantum dynamics. A natural question arises: can Theorem 1.4 be complemented to yield such a characterization? As noted in Remark 1.6, the sequences of surfaces may in fact possess slowly decaying injectivity radii and spectral gaps. Given the connection between weak mixing and the spectral gap, it is plausible to conjecture that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum weak mixing if and only if the spectral gap tends to zero sufficiently slowly. Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable.
While full characterization results may still be out of reach, the present results appear extendable—albeit with varying levels of difficulty. Several directions for possible extensions present themselves. For instance, can analogous results be established in other structures such as graphs? We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal elements by adapting the techniques developed here, thereby enabling a proof of large-scale weak mixing in the graph setting of Anantharaman and Le Masson. Doing quantum mixing on the graphs seems highly timely as Bordenave, Letroit, and Sabri have an upcoming result on mixing on growing Schreier graphs [BLS].
Another avenue of extensions concerns the choice of observables: in Theorem 1.1, the observable class was quite broad—zeroth-order pseudodifferential operators—so it seems likely that with minor modifications, a larger class of observables could be accommodated in the large-scale setting as well. This has already been achieved in the large-scale quantum ergodicity contexts for graphs [AL15] and surfaces [ABL23], where the kernels of the observables are concentrated near the diagonal but not restricted to it, unlike in the present work.
Yet another natural question is whether the theorem can be adapted to the case of variable curvature, or to accommodate more general energy operators than the Laplacian. For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with general Schrödinger operators, under suitable restrictions. Non-regular graphs can be viewed as discrete analogues of variable-curvature surfaces. It would also be of great interest to construct counterexamples—sequences of surfaces demonstrating that the conditions imposed on the sequences are indeed necessary. In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by MacKenzie [McK22]; however, constructing such examples in the surface setting remains substantially more difficult.
We pose perhaps an obvious question: is it possible to remove the averaging present in our theorems? In the large-energy regime, this corresponds to the quantum unique ergodicity conjecture of Rudnick and Sarnak, see [RS94]. The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces [Lin06]), but remains open in general. Analogous large-scale results are known for graphs; for instance, large-scale quantum unique ergodicity for random regular graphs was proved by Bauerschmidt et al. [BHY19]. We conjecture that the conclusions of Theorem 1.4 remain valid even without averaging. We refer to such statements as large-scale quantum unique ergodic and large-scale quantum unique mixing. These would parallel the averaging-free results of Cipolloni et al. [CES21]. A possibility is to try to establish this first using random models, such as Weil-Petersson model, or by using arithmetic surfaces and thus complement Lindenstrauss [Lin06]. In particular, for arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected via the Watson $L$-function [Wat02].
In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos—such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [Deu18]. Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance of the eigenvalues of $\psi_{j}$ and $\psi_{k}$ goes to infinity. Looking at the large-scale results discussed here, Theorems 1.4 and 1.7, we see that parameter $\tau$ controls the distance of the eigenvalues. However, the eigenvalues are taken from the fixed energy windows which means that for large enough $\tau$, the averages of the off-diagonal elements will be zero as the large distance is not possible in the energy window. However, if we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions. It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will close to zero, but verifying this requires careful study of the sequence of growinf surfaces, $\tau$, and the energy window.
2. Preliminaries
In this section, we recall some basic concepts and results that will be used throughout this article, without proofs. For background on hyperbolic geometry concerning the hyperbolic plane and hyperbolic surfaces, we refer to Katok [Kat92], and for details on the Laplace–Beltrami operator, to Iwaniec [Iwa02].
2.1. The hyperbolic plane and hyperbolic surfaces
The hyperbolic plane $\mathbb{H}$ is identified with the upper half-plane
$\{x+iy\in\mathbb{C}:y>0\},$
equipped with the Riemannian metric
$ds^{2}=y^{-2}(dx^{2}+dy^{2}).$
For points $z,z^{\prime}\in\mathbb{H}$, we denote by $d(z,z^{\prime})$ the hyperbolic distance between them. The associated volume measure is
$d\mu(x+iy)=y^{-2}\,dx\,dy,$
and we write $|A|$ for the hyperbolic volume of a set $A\subseteq\mathbb{H}$.
The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm I$, acts on $\mathbb{H}$ by Möbius transformations:
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot z=\frac{az+b}{cz+d}.$
This action identifies $PSL_{2}(\mathbb{R})$ with the group of orientation-preserving isometries of $\mathbb{H}$.
The unit tangent bundle of $\mathbb{H}$ is
$T^{1}\mathbb{H}=\{(z,v)\in\mathbb{H}\times T_{z}\mathbb{H}:\|v\|_{z}=1\}.$
The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot(z,v)=\left(\frac{az+b}{cz+d},\frac{v}{(cz+d)^{2}}\right).$
Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$; hence, these two spaces can be identified.
The geodesic flow is the one-parameter family of maps
$\varphi_{t}:T^{1}\mathbb{H}\to T^{1}\mathbb{H},\qquad t\in\mathbb{R},$
such that for any $(z,v)\in T^{1}\mathbb{H}$, the curve $(\eta(t),\eta^{\prime}(t))=\varphi_{t}(z,v)$ is the unique geodesic parametrized by arc length satisfying $\eta(0)=z$ and $\eta^{\prime}(0)=v$. The geodesic
$(i,i)\mapsto\varphi_{t}(i,i)=\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}\cdot(i,i)$
corresponds to the imaginary axis. For any $(z,v)\in T^{1}\mathbb{H}$, there exists $g\in PSL_{2}(\mathbb{R})$ such that $g\cdot(i,i)=(z,v)$, and thus
$\varphi_{t}(z,v)=g\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}\cdot(i,i).$
Hence, identifying $T^{1}\mathbb{H}$ with $PSL_{2}(\mathbb{R})$, the geodesic flow corresponds to the right action of the one-parameter subgroup
$A=\left\{a_{t}=\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}:t\in\mathbb{R}\right\}.$
We denote by $z,z^{\prime}]$ the geodesic segment connecting two distinct points $z,z^{\prime}\in\mathbb{H}$, by $z,z^{\prime}](w)$ the tangent direction at a point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ its length.
For any distinct points $z,z^{\prime}\in\mathbb{H}$, there exists a unique geodesic of length $r$ from $z$ to $z^{\prime}$. Using the geodesic flow, there exist unique $\theta\in\mathbb{S}^{1}$ and $r\in(0,\infty)$ such that $z^{\prime}$ is the projection of $\varphi_{r}(z,\theta)$ onto the first coordinate. The change of variables $z^{\prime}\mapsto(r,\theta)$ defines the polar coordinates around $z$, in which the metric and volume element are
$ds^{2}=dr^{2}+\sinh^{2}(r)\,d\theta^{2},\qquad d\mu(r,\theta)=\sinh(r)\,dr\,d\theta.$
Then a point $x\in\mathbb{H}$ can be written as
$x=\pi_{1}(\phi_{r}(z,\theta)),$
where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of pair in $T^{1}\mathbb{H}$.
The Laplacian on $\mathbb{H}$ is the differential operator
$\Delta=-y^{2}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right),$
which commutes with all isometries of $\mathbb{H}$.
2.2. Hyperbolic surfaces
A compact connected hyperbolic surface can be realized as a quotient
$X=\Gamma\setminus\mathbb{H},$
where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$. In this way, $X$ can be represented by a fundamental domain in $\mathbb{H}$.
A function on $X$ can equivalently be regarded as a $\Gamma$-periodic function on $\mathbb{H}$, and vice versa. Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain.
For a hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, the unit tangent bundle can be identified with $\Gamma\setminus PSL_{2}(\mathbb{R})$, and under this identification the geodesic flow takes the form
$\varphi_{t}(\Gamma g)=\Gamma g\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}$.
The injectivity radius of $X$ at a point $z\in X$ is defined by
$\operatorname{InjRad}_{X}(z):=\tfrac{1}{2}\min_{\gamma\in\Gamma\setminus\{\mathrm{id}\}}d(z,\gamma z),$
and the (global) injectivity radius of $X$ is
$\operatorname{InjRad}_{X}:=\sup_{z\in X}\operatorname{InjRad}_{X}(z).$
Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally to a Laplacian on the quotient $X=\Gamma\setminus\mathbb{H}$. To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on all smooth functions on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions. On a compact connected hyperbolic surface, the Laplacian $\Delta_{X}$ has a discrete spectrum
$0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\dotsb\xrightarrow{j\to\infty}\infty.$
3. Strategy of proofs
Three main theorems, Theorems 1.4, 1.5, and 1.7, are proven in Section 8. In this section we give a rough, overarching idea of the approach taken in the proofs. Most of the paper is dedicated for trying to establish Theorem 1.5. Proving this theorem is technical so the proof is split into multiple sections. These sections, Sections 4, 5, 6, and 7, are presented before the Section 8 so that the auxiliary results have been covered when the main theorems are under focus.
In Theorem 1.5 we want to bound the double sum appearing in Expression (3.3) for a fixed surface $X=\Gamma\setminus\mathbb{H}$ under suitable conditions. Hence we want to bound the following:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}$ (3.1)
We can split the inner product as follows:
$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.2)
when
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t}(x)=\frac{\sin(tx)}{x},$
and
$P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t,\tau}=\cos(t\tau)\,h_{t}.$
This split could be referred to as geometric-spectral split, since the denumerator contains the spectral data of the problem and the numerator the geometric data of the problem. $P_{t}$ is the hyperbolic wave propagator (see [LP82]). It is the quantization of the geodesic flow, hence a very natural object. Using this propagator makes the spectralside easy but complicates the analysis of the geometric side when comparing the setting to that of Le Masson and Sahlsten (see [LS17]) who used a ball averaging propagator which is not as natural as the wave propagator. $P_{t,\tau}$, a modified version of $P_{t}$, is used simultaneously with $P_{t}$ so that we can access the off-diagonal elements. This adds another layer of complexity to the proof when comparing to the setting in which only the diagonal elements are considered.
The denumerator of Expression (3.2) contains the spectral data of the problem. We acquire bounds for the denominator in Section 4. Using the bounds acquired there gives an upper bound for Expression (3.1):
$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2},$
where the Hilbert-Schmidt norm upper bound was acquired by enlargening the sums to include all the basis elements after the spectral bound was pulled out of the sum. Now the Hilbert-Schmidt norm contains the geometric data of the problem.
Bouding the Hilbert-Schmidt norm is very technical. In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$. Using the acquired formulas, we bound the Hilbert-Schmidt norm in Section 6. This is based on carefully bounding the Hilbert-Schmidt norm expression and finally splitting it into two cases. In one we consider an integral over points with small injectivity radius and in the other an integral over points with large injectivity radius. To bound these two cases, we especially need the exponential mixing theorem, Theorem 1.2, and bounds for certain geometric weight $F_{t,t^{\prime},\rho}$ which appears in both the small and large injectivity radius cases, containing the core geometric information of the problem. Bounding $F_{t,t^{\prime},\rho}$ is done in Section 7. We remark that the exponential mixing theorem replaces the Nevo’s theorem appearing in the same role in [LS17] as the exponential mixing here.
With the Hilbert-Schmidt norm bounded, we acquire the final bounds for Expression (3.1) to conclude Theorem 1.5:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq E\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we refer to Theorem 1.5 for the quantities appearing in the upper bound.
The limit version of the main theorems, Theorem 1.4, is proven using Theorem 1.5 together with the large-scale Weyl law of Le Masson and Sahlsten, Theorem 8.1. In the proof we want to prove that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have that
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}\right)<\epsilon.$ (3.3)
Proving this proves only conditions (ii) and (iii); Le Masson and Sahlsten’s proof of quantum ergodicity covers case (i), see [LS17]. We remark here that the techniques used here can also be used to give a new proof for the quantum ergodicity theorem. Using the bounds acquired for the double sum in Theorem 1.5 gives the following upper bound for the previous expression:
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,E\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}(X_{n}))^{3}}\,\left(\|a_{n}\|_{2}^{2}\,+\,8\pi^{2}\|a_{n}\|_{\infty}^{2}\,|D_{n}\setminus D_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right)\right),$
where we refer to Theorem 1.4 for the quantities appearing in the upper bound. Using property (BSC) helps us to get a universal lower bound for $\beta(\lambda_{1}(X_{n}))$ and using property (UND) helps us to get a universal lower bound for $\operatorname{InjRad}_{X_{n}}$. Weyl law together with property (BSC) shows that
$\frac{|D_{n}\setminus D_{n}(4T)|}{|N(X_{n},I)|}\xrightarrow{n\to\infty}0,$
as the Weyl law relates the growth of the number of eigenvalues in the energy window $I$ to that of the volume of the surface, $|X_{n}|$, and the ratio of the points with small injectivity radius and the volume of the surface goes to zero due to property (BSC). The core idea is to use properties (BSC), (EXP), and (UND) together with Weyl law to remove the limit supremum from the upper bound. After this, we have an upper bound in which $\delta$ appears. We can choose a suitable $\delta$ to get $\epsilon$ as an upper bound to conclude the proof.
To finally prove the probabilistic main theorem, Theorem 1.7, we take a very similar approach as with the proof of Theorem 1.4 as these theorems are very similar. We only prove Properties (ii) and (iii) of Theorem 1.4 as Le Masson and Sahlsten have already proved Property (i), see [LS24]. We remark that in probabilistic setting we do not impose Properties (BSC), (EXP), and (UND) and hence we do not have the large-scale Weyl law at out disposal. To remedy this, we get similar properties from the theory of random surfaces. These are recorded in Theorem 8.2. We can use these properties in the place of Properties (BSC), (EXP), and (UND) in a very similar way to get the conclusion to hold with high probability as we proceed in the sequence.
To conclude the section, we remark that the main novelty of this paper method-wise is the use of the exponential mixing theorem, Theorem 1.2, instead of using Nevos’s theorem to get the needed extra decay to establish Theorem 1.5. Another novelty is to separate a weight arising from the choice of propagator, denoted here by $F_{t,t^{\prime},\rho}$, which gives much required nimbleness in the proof; it can also be approximated separately without the observable function $a$ interfering.
4. Spectral data
In this section we prove a bound for spectral quantity
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}$
under certains assumptions on $a,b,t,\tau,$ and $T$, given that
$h_{t}(x)=x^{-1}\,\sin(tx),\qquad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
The spectral quantity appears in the proof of Theorem 1.5 and there it contains the spectral data of the problem.
We state the proposition.
Proposition 4.1 (Spectral data).
Let $\tau\in\mathbb{R}$. Assume that $0<m<a,b<\infty$ and that
$a-b-\tau\in(-\delta,\delta),$
where $0<\delta<\tfrac{2}{9}\,m$ and where $T\,\delta=\pi/2$. Then
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}\,<\,8\pi ab.$
The proof consists of relatively straightforward integration and careful application of the relationships of the variables and coefficients. The simplicity follows from a convenient choice of $h_{t}$ and $h_{t,\tau}$.
Proof of Proposition 4.1.
We set out to bound
$G_{\tau,T}=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|$
from below to prove the bound. $G_{t,\tau}$ is illustrated in Figure 1.
Figure 1. Contour plot of $G_{\tau,T}(a,b)$. Two pronounced “ridges” are visible along the lines $b=a+\tau$ and $b=a-\tau$. The ridges get pronounced as $T$ increases. For a fixed $\delta$, $G_{\tau,t}(a,b)$ has a positive lower bound when $a-b-\tau\in(-\delta,\delta)$. Notice also how the ridges decay as $a$ and $b$ grow.
By using the definitions of $h_{t}$ and $h_{t,\tau}$, we write $G_{\tau,T}(a,b)$ as
$\frac{1}{Tab}\left|\int_{0}^{T}\cos(t\tau)\sin(ta)\sin(tb)dt\right|.$ (4.1)
By using trigonometric sum-to-product identities, we have that $\cos(t\tau)\sin(ta)$ equals
$\frac{1}{2}\left(\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)\right),$
and further that $\cos(t\tau)\sin(ta)\sin(tb)$ equals
$\frac{1}{4}\left(\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\right).$
Using this equality, we have that Expression (4.1) equals
$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\,dt\right|.$
The sine terms of the integral can all be integrated separately and thus the integral yields the following equal expression:
$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|$
Applying the reverse triangle inequality yields a following lower bound for the previous expression:
$\frac{1}{4Tab}\bigg(\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|\bigg).$ (4.2)
We proceed to bound the absolute value terms separately. Consider the first absolute value term of the previous expression. Consider function
$x\mapsto\frac{\sin(T\,x)}{x}.$
It is positive when $x\in\left(-\pi/T,\pi/T\right)$, and it also decreases as $|x|$ grows in this interval. Because $T\delta=\pi/2$ and $a-b-\tau\in(-\delta,\delta)$, we can apply the observations regarding the previously studied function to get that
$\frac{\sin\!\big(T(a-b-\tau)\big)}{a-b-\tau}\ \geq\ \frac{\sin(T\delta)}{\delta}=\frac{1}{\delta}.$
We also notice that
$\frac{\sin(Tx)}{x}>0$
when $x\in\left(-\pi/T,\pi/T\right)$, which means that for every $x\in\mathbb{R}$ we have that
$\frac{\sin(Tx)}{x}\geq-\frac{1}{\pi/T}=-\frac{1}{2\delta}.$
The previous two observations mean that
$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}$
is necessarily positive and hence
$\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|=\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
Next, consider the second and third absolute value terms of Expression (4.2). We observe that
$\left|\frac{\sin(Tx)}{x}\right|\leq\frac{1}{|x|},$
implying that
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}.$
We also observe that
$|a+b-\tau|\geq 2b-|a-b-\tau|>2m-\delta>8\delta,$
where we used the reverse triangle inequality together with the fact that $|a-b-\tau|$ is smaller than $\delta$ and that $2m$ is larger than $9\delta$. Using these observations gives that
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}<\frac{1}{8\delta}.$
A similar approach yields that
$\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta}.$
These two observations mean that
$-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|>-\frac{1}{4\delta}.$
With all the absolute value terms bounded, Expression (4.2) has the following lower bound:
$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab}.$
In the previous we used that $T\delta=\pi/2$.
Combining these bounds gives that
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|\ \geq\ \frac{1}{4Tab}\cdot\frac{1}{4\delta}\ =\ \frac{1}{16\,\frac{\pi}{2}\,ab}\ =\ \frac{1}{8\pi\,ab}.$
Therefore,
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab,$
as claimed. ∎
5. Propagators
In order to estimate the Hilbert–Schmidt norm
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly mofified version of it, $P_{t,\tau}$, for these kernels are then used to open the Hilbert-Schmidt norms. Although the hyperbolic wave propagator is a natural choice, it exhibits greater thechnical complexity than, for instance, the ball averaging operator (cf. [LS17]). As mentioned earlier, this choice makes the specral side easier.
We state a proposition giving a convenient form for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
Proposition 5.1 (Propagators).
Let
$P_{t}=h_{t}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),\quad P_{t,\tau}=h_{t,\tau}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),$
be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with $\Delta_{X}$ being the Laplace-Beltrami operator and where
$h_{t}(x)=\frac{\sin(tx)}{x},\quad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
Then $P_{t}$ admits integral kernel
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
and $P_{t,\tau}$ admits integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
Given that
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right)$
and that $\tilde{P}_{t}$ is the operator with the integral kernel $K_{t}^{\Gamma}$ which was presented in the statement of the proposition, the proposition can be proved by showing that inner product $\langle P_{t}f,g\rangle$ equals inner product $\langle\tilde{P}_{t}f,g\rangle$ where $f,g\in C^{\infty}(X)$ as $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ with $X$ being compact. The core idea is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have a convenient explicit formulae for their integral kernels.
The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1 for clarity’s sake
Proof of Proposition 5.1.
The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence it suffices to consider $P_{t}$ and $K_{t}$.
Due to way $P_{t}$ is defined, it is bounded and has property
$P_{t}\psi_{j}=h_{t}(\rho_{j})\,\psi_{j},$
where $\psi_{j}$ is an eigenfunction of the Laplacian on $X$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-1/4}$. Define regularized operators for $\varepsilon>0$:
$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
Due to way $P_{t}^{(\epsilon)}$ is defined, it is bounded and has property
$P_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\,\psi_{j}.$
Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon$ shrinks, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise as $\varepsilon$ shrinks for every $f\in C^{\infty}(X)$.
Proposition 5.2 ensures the existence of integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel
$K_{t}^{(\varepsilon)}(x,y)=\sum_{\gamma\in\Gamma}k_{t}^{(\varepsilon)}(x,\gamma y),$
$k_{t}^{(\varepsilon)}(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\Bigg(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\Bigg)\rho h^{(\varepsilon)}(\rho)\,d\rho,$
such that
$Q_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\psi_{j}.$
Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form so that
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}$
denotes the standard Gaussian mollifier. Since an operator on $X$ is uniquely determined by its action on the orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
Let $\tilde{P}_{t}$ denote an integral operator on $X$ with kernel $K_{t}^{\Gamma}$. To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS12]), so it suffices to verify that for all $x\in X$
$\int_{X}K_{t}(x,y)\,dy$
is finite as $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. Open the previous expression using the definition of $K_{t}$:
$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\bm{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
Uniting the integral and the sum into an integral over the universal cover gives the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
Writing this using polar coordinates around $x$ yields
$\frac{2\pi}{2\sqrt{2}\pi}\int_{0}^{t}\frac{\sinh(\rho)}{\sqrt{\cosh(t)-\cosh(\rho)}}d\rho=\frac{4\pi}{2\sqrt{2}\pi}\sqrt{\cosh(t)-1},$
where coefficient $2\pi$ appeared first from the polar coordinates since there was no dependency on the direction, and the equality followed from computing the integral. The result is finite and $\tilde{P}$ is well-defined and bounded.
To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the statement of the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$
$\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle=0\iff\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, start by considering the absolute value term of the previous expression. It can be written as
$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|,$
where we opened the inner product and united the two integrals. By pulling the absolute values inside the integral and then bounding $|g|$ with the $\sup$ norm, we get the following upper bound for the previous expression:
$\|g\|_{\infty}\int_{X}\left|P_{t}f(x)-\tilde{P}_{t}f(x)\right|dx.$
Focus on the absolute value integrand appearing in the previous expression. We can write it in the following form by writing $P_{t}f$ using the limit and then by writing the operator action using the kernels:
$\left|\lim_{\varepsilon\to 0}\int_{X}K_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}^{\Gamma}(x,y)f(y)dy\right|.$
Writing $K_{t}^{\Gamma}$ and $K_{t}^{(\varepsilon)}$ using their sum representations over $\Gamma$ and then uniting the integral and the sum into an integral over the universal cover yields the following equal expression, where we also view $f$ as $\Gamma$ periodic function:
$\left|\lim_{\varepsilon\to 0}\int_{\mathbb{H}}k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)f(y)dy\right|.$
By pulling the absolute values inside the integral and bouding $|f|$ with the $\sup$ norm, we get the following upper bound:
$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)\right|dy.$
Focus on the integral in the previous expression. By opening $k_{t}$ and $K_{t}$, the integral can be written as
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{\left(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)\right)\,\bm{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
By defining
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
we may write the preceding expression as
$\frac{1}{2\sqrt{2}\pi}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(t)+\chi_{d(x,y)}(-t)\right|\,dy,$
where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$. The additional term $\chi_{d(x,y)}(-t)$ equals 0 so adding it had no effect on the equality. Using the triangle inequality, the previous expression is bounded by
$\frac{1}{2\sqrt{\pi}}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|\,dy.$
Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. The line of reasoning leading to this cocnlusion means that
$\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle,$
thus concluding the proof. ∎
The remaining section contains the lemmas and propositions used in the proof of Proposition 5.1. First, we record a slightly modified formulation of Proposition 5 in [Mar04]. It gives a family of operators with explicit formulae for their integral kernels on the compact connected hyperbolic surface. Especially, these operators share the eigenfunctions with the Laplacian of the surface and of which eigenvalues have a representation using the eigenvalues of the Laplacian.
Proposition 5.2.
Let $h:\{x+iy:|y|<C\}\to\mathbb{C}$ satisfy:
  1. i.
    $h$ is analytic,
  2. ii.
    $h$ is even,
  3. iii.
    $h(x+iy)$ decays superpolynomially as $|x|\to\infty$,
[…]
$K^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}K(x,\gamma y),$
where
$K(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\left(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\right)\rho h(\rho)\,d\rho,$
satisfies property
$P\psi_{j}=h(\rho_{j})\psi_{j},$
where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-1/4}$.
Next, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, an integral kernel that is crucial in establishing a convenient formula for the integral operator $P_{t}$.
Lemma 5.3.
Let
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{\pi i}\int_{-\infty}^{\infty}\left(-\frac{1}{2\pi\sqrt{2}}\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\right)\rho h_{t}^{(\varepsilon)}(\rho)\,d\rho,$
where $x,y\in\mathbb{H}$ and
$h_{t}^{(\varepsilon)}(\rho)=\frac{\sin(t\rho)}{\rho}\,e^{-\frac{\varepsilon^{2}\rho^{2}}{2}}.$
Then
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}$
is the standard Gaussian mollifier.
The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$ in the statement. Then by using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(t\,x)$ (in the distributional sense) together with the Fourier convolution identity, we acquire the wanted conclusion.
Proof of Lemma 5.3.
Applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
$\frac{1}{\pi i}\frac{1}{2\sqrt{2}\,\pi}\int_{d(x,y)}^{\infty}\frac{1}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\left(\int_{-\infty}^{\infty}e^{-i\rho v}\rho h_{t}^{(\varepsilon)}(\rho)\,d\rho\right)dv.$ (5.1)
The inner integral is the Fourier transform of $\rho\mapsto\rho h_{t}^{(\varepsilon)}(\rho)$, where the Fourier transform is
$\mathcal{F}(f)(x)=\int_{\mathbb{R}}f(y)e^{-ixy}\,dy$
and hence the Inverse Fourier transform is
$F^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
We observe that
$\rho\,h_{t}^{(\varepsilon)}(\rho)=\sin(t\rho)\,e^{-\frac{\varepsilon^{2}\rho^{2}}{2}}=2\pi\,\sin(t\rho)\,\mathcal{F}^{-1}(\phi_{\varepsilon})(\rho),$
since
$\mathcal{F}^{-1}(\phi_{\varepsilon})(x)=\frac{1}{2\pi}\,e^{-\frac{\varepsilon^{2}\,x^{2}}{2}}.$
By using the Fourier convolution identity and the Fourier transform of sine we observe that the inner integral of Expression (5.1) equals
$\mathcal{F}\left(\rho\mapsto 2\pi\,\sin(t\rho)\,\mathcal{F}^{-1}(\phi_{\varepsilon})\right)(v)=\frac{2\pi}{2\pi}\left(\mathcal{F}(s_{t})\ast\phi_{\varepsilon}\right)(v)=\pi i\left(\phi_{\varepsilon}(v+t)-\phi_{\varepsilon}(v-t)\right),$
where $s_{t}$ denotes function $x\mapsto\sin(tx)$. Substituting this into Expression (5.1) yields the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where the equality followed using the symmetry of the Gaussian mollifier.
Finally, the following lemma establishes convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Interestingly, the convolution here is over the real line but the convergence is studied in the $\mathbb{H}$.
Lemma 5.4.
Let
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}$
and let
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}.$
Then map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$.
In the proof we show that
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy\xrightarrow{\varepsilon\to 0}0.$
The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)$, where $A$ is a continuous function with $A(t)=0$. Then the conclusion can be drawn using the properties of the convolution.
Proof of Lemma 5.4.
We want to show that the following expression goes to 0 as $\varepsilon$ goes to zero:
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$
By the triangle inequality and Fubini’s theorem, this is bounded by
$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A(t),$
where
$A(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
By using the polar coordinates around $x$ together with the triangle inequality we have that
$A(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
Hence we observe that $v\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function with which the dominated convergence theorem can be used to observe that $A(v)\to 0$ as $v\to t$. $A$ is continuous, and as it is continuous, we have that
$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)=0,$
establishing the claim. ∎
6. Geometric data
In this section, we bound the Hilbert-Schmidt norm
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
which is the geometric data appearing in the proof of Theorem1.5 in Section 8. $P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5.
Proposition 6.1 (Geometric data).
Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X$ with Laplace-Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Then
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D_{1}\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
where $\beta=\beta(\lambda_{1}(X))$, where $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{7}=\frac{7200}{\pi}.$
This is the most technical part of the whole paper. The core idea is first to write $K_{t}^{\Gamma}$ and $K_{t,\tau}^{\Gamma}$ using $K_{t}$ while simultaneously simplifying the expressions by unifying the sums over the Fuchsian group $\Gamma$ with the integrals over the fundamental domains of the surface $X$ to get integrals over the whole hyperbolic plane. This careful splitting allows us to separate a certain weight function $F_{t,t^{\prime},\rho}$ into a separate integral in which it is easier to bound. Section 7 is dedicated for the analysis of $F_{t,t^{\prime},\rho}$. Finally, we split the expression into two sides, one with the points with large injectivity radius and the other with the points with small injectivity radius. Both cases require careful analysis to acquire reasonable bounds, especially the side with the large injectivity radius for which we need the exponential mixing theorem, Theorem 1.2.
The proof relies on multiple auxiliary lemmas presented after the proof. This approach makes the proof clearer.
Proof of Proposition 6.1.
Writing the Hilbert-Schmidt norm as a double integral over the fundamental domain $D$ of $X$ yields the following expression:
$\int_{D}\int_{D}\left|\frac{1}{T}\int_{0}^{T}\int_{D}K_{t,\tau}^{\Gamma}a(z)K_{t}^{\Gamma}(z,y)dt\right|^{2}dydx.$
Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
$\int_{D\times D}\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dz^{\prime}dzdt^{\prime}dtdydx.$
Using Fubini’s theorem to reorder the integrals yields the following equals expression:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D\times D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dxdydz^{\prime}dzdt^{\prime}dt.$
This expression is bounded by adding the absolute values around it. Then we pull the absolute values inside to get the following upper bound:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}\int_{D}\int_{D}K_{t}^{\Gamma}(x,z)a(z)K^{\Gamma}_{t}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dydxdz^{\prime}dz\right|dt^{\prime}dt.$
This can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same thus resulting into an integral squared:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{D}K_{t}^{\Gamma}(x,z)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt.$
Lemma 6.2 allows us to write the previous expression using the hyperbolic plane kernels $K_{t}$ and $K_{t^{\prime}}$ instead of the surface kernels $K_{t}^{\Gamma}$ and $K_{t^{\prime}}^{\Gamma}$:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt$
Coefficient $(8\,\pi^{2})^{-1}$ came from the definition of $K_{t}^{\Gamma}$:
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y).$
Lemma 6.3 allows us to rewrite the integral inside the absolute value of the previous integral expression in a very convenient form, eliminating but one sum. Thus the previous expression equals:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{\mathbb{H}}a(z)a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K_{t}(x,z)K_{t^{\prime}}(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K_{t}(y,\gamma z)K_{t^{\prime}}(y,z^{\prime})dy\right)dz^{\prime}dz\right|dt^{\prime}dt$
which further equals
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
when $F_{t,t^{\prime},\rho}$ is defined to equal
$\int_{\mathbb{H}}K_{t}(x,w)K_{t^{\prime}}(x,w^{\prime})dx,$
where $w,w^{\prime}\in\mathbb{H}$ are any points seprated by distance $\rho$. In Section 7, where we study $F_{t,t^{\prime},\rho}$ carefully, in Lemma 7.1 we will see that the chosen points $w,w^{\prime}$ do not matter, only their distance.
The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, which follows from the definition of $K_{t}$ and $K_{t^{\prime}}$. Using this observation gives us the following equal expression for Expression (6.1):
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
Focus on the absolute values in the previous expression. They are bounded from above by
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$ (6.3)
where the integration over $D$ is split into two: the points having injectivity radius bigger than $4T$, $D(4T)$, and the points having injecitivity radius smaller than $4T$, $D\setminus D(4T)$. The large injectivity radius case and the small injectivity radius case are bounded separately.
The small injectivity radius term is bounded in Lemma 6.4. Using its bound gives the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
In the previous, the second term is bounded by the third term as the second term has no sum and the third one does. Hence the previous expression is bounded by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
+ $+\ 2\,\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$
where the second term is the small injectivity radius case, which is bounded in Lemma 6.5. Using this lemma gives the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$ (6.4)
The latter term is bounded by
$8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}11e^{\beta}\sinh(\rho)(1+\rho)e^{-\beta\rho}e^{2\beta T}F_{t,t^{\prime},\rho}^{2}d\rho,$
since
$11\,e^{\beta}(1+\rho)e^{-\beta\rho}e^{\beta\,2T}\geq 1.$
Hence Expression (6.4) is bounded by:
$11e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
We have now acquired good bounds for the absolute values appearing in Expression (6.1), and this expression is thus bounded by
$\frac{11}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
As $F_{t,t^{\prime},\rho}$ equals $F_{t^{\prime},t,\rho}$ and since we integrate from 0 to $T$ both $t$ and $t^{\prime}$, we can use symmetry to integrate the inner integral of $t^{\prime}$ from 0 to only $t$ twice. Hence we write the previous expression as
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$ (6.5)
The innermost integral can be bounded using Proposition 6.6, which uses the bound acquired for $F_{t,t^{\prime},\rho}$ in Section 7. We get the following upper bound for the previous expression:
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}dt^{\prime}dt,$
where
$C_{6}=480\pi.$
The remaining two integrals have
$\frac{2T\,C_{6}}{\beta^{3}}$
as an upper bound, which is seen after evaluating these integrals. Then the previous integral expression has the following upper bound:
$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$\frac{22\cdot 2\cdot C_{6}\cdot e}{8\pi^{2}}\leq\frac{7200}{\pi}=C_{7},$
where we used that $e^{\beta}$ is less than $e$. This is also an upper bound for the Hilbert-Schmidt norm we set out to bound. It is of the wanted form which concludes the proof.
The remaining of this section contains the lemmas needed to prove Proposition 6.1 together with the contents of Section 7. The following lemma allows us to write the integration containing the kernels living on the hyperbolic surface using the kernels living on the hyperbolic plane.
Lemma 6.2.
Let
$K^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}K(x,\gamma y),\qquad L^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}L(x,\gamma y)$
be integral kernels on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, where $K$ and $L$ are integral kernels on $\mathbb{H}$ that depend only on distance $d(x,y)$. Then
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma_{2}z^{\prime})dx,$
where $D$ is a fundamental domain of $X$.
The proof is on uniting the sum over the Fuchsian group and the integration over the fundamental domain into an integration over the hyperbolic plane.
Proof of Lemma 6.2.
We have that
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{2}z^{\prime})\,dx$
by opening the kernels using their sum form. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{1}\gamma_{2}z^{\prime})dx.$
Using that $K(x,\gamma_{1}z)=K(\gamma_{1}^{-1}x,z)$ and that $L(x,\gamma_{1}\gamma_{2}z^{\prime})=L(\gamma_{1}^{-1}x,\gamma_{2}z^{\prime})$ allows us then to write the previous expression as:
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(\gamma_{1}^{-1}x,z)L(\gamma_{1}^{-1}x,\gamma_{2}z^{\prime})dx.$
Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})\,dx,$
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof.
The following lemma allows us to simplify the notation by removing one sum over the Fuchsian group.
Lemma 6.3.
Let $K$ and $L$ be integral kernels on compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$. Let $a$ be a function on $X$. Then:
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
$=\qquad$ $\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
where $D$ is a fundamental domain of $X$.
The proof is based on the same principle as in the proof of Lemma 6.2: we unite the sum over the Fuchsian group and the integral over the fundamental domain into an integration over the hyperbolic plane.
Proof of Lemma 6.3.
Focus on the following expression:
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
We expand the squares and reorder the integrals and the sums using Fubini’s theorem to get the following expression:
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
which further equals
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
when using that $L(x,\gamma y)=L(\gamma^{-1}x,y)$. Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as $\Gamma$-periodic function on $\mathbb{H}$. This gives the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.6)
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. For the latter integral we have that
$\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy=\int_{\mathbb{H}}K(\gamma^{-1}y,z)L(y,z^{\prime})dy=\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy,$
where the first equality follows from the change of variables $y\mapsto\gamma^{-1}y$, and the second equality follows from the fact that $L(\gamma^{-1}x,y)=L(x,\gamma y)$. With this observation, we can rewrite Expression (6.6) as
$\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
The following lemma deals with the integration over the part of the fundamental domain with large injectivity radius, giving us the upper bound consisting of two terms: one is an integral with integrand with exponential decay and the other is an integration over the points with the small injectivity radius. The first term is desired and can be united with the bound acquired from the small injectivity radius case, but the second term needs to be further bounded. This is taken care of when we deal with all the parts dealing with the parts with small injectivity radius, which is done in Lemma 6.5.
Lemma 6.4.
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius bigger than $4T$. Let $a$ be a mean zero function on $X$. Then
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right|$
$=\ \|a\|_{2}^{2}\,\int_{0}^{2T}\,11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}\,(1+\rho)\,e^{-\beta\rho}\,d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|,$ (6.7)
where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of Laplace-Beltrami operator $\Delta_{X}$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The proof is based on modifying the expression so that the exponential mixing theorem, Theorem 1.2, can be applied.
Proof of Lemma 6.4.
As $z\in D(4T)$, we have that $d(z,\gamma z)$ is larger than $4T$ for any non-identity group element $\gamma$. We also have that $z^{\prime}\in B(z,2T)$. This means that $d(\gamma z,z^{\prime})$ is larger than $2T$ for any non-identity group element. As $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, we have that only the identity element can have a non-zero contribution. Expression (6.4) equals
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.8)
By adding and subtracting
$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
in the absolute values of Expression (6.8) and then appyling the triangle inequality gives the following expression as an upper bound for Expression (6.8):
$\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.9)
Focus on the first term. We write $z^{\prime}$ using polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, and the first term equals
$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a^{*}(\varphi_{\rho}(z,\theta))\,F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz\right|,$
where $a$ can be seen as a function on the unit tangent bundle by defining that $a(z,\theta)=a(z)$. The integrals of this expression can be reordered using Fubini’s theorem yielding the following expression:
$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a^{*}(\varphi_{\rho}(z,\theta))\,d\theta\,dz\,d\rho\right|.$
By pulling the absolute values inside the first integral, we have the following upper bound, where the inner two integrals are written as an inner product:
$\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,\left|\left\langle a,a\circ\varphi_{\rho}\right\rangle_{L^{2}(D\times\mathbb{S}^{1})}\right|d\rho.$
We can apply the Exponential mixing theorem, Theorem 1.2, to the absolute value of the inner product which then yields the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho.$
Hence Expression (6.9) and hence the original expression we set out to bound are bound by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$
The following lemma deals with the integration over the part of the fundamental domain with small injectivity radius giving an upper bound for it. This can then be united with the bound acquired for the large injectivity radius case in Lemma 6.4.
Lemma 6.5.
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius larger than $4T$. Then
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
is bounded from above by
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho$
The proof is based on careful use of Hölder’s inequality together with a careful management of the sum elements with non-zero contribtution.
Proof of Lemma 6.5.
Consider
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
By pulling the absolute values inside this is bounded by
$\int_{D\setminus D(4T)}|a(z)|\int_{B(z,2T)}\left|a^{*}(z^{\prime})\right|F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$
which is further bounded by the following expression when using the supremum norm bounds for $|a|$ and $|a^{*}|$:
$\|a\|_{\infty}^{2}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
In the previous expression, the sum elements with a possible non-zero contribution are those $\gamma\in\Gamma$ such that $d(\gamma z,z^{\prime})$ is smaller than $2T$ for some $z^{\prime}\in B(z,2T)$. This collection depends on $z$ and we denote it by $\Gamma(z)$. We have that
$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}},$
which follows from that the minimum distance of $\gamma z$ and $\gamma^{\prime}z$, where $\gamma\neq\gamma^{\prime}$ is $\operatorname{InjRad}_{X}$, and that $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})$ is smaller than $2T$, since $z^{\prime}\in B(z,2T)$. Hence the previous integral expression has the following upper bound when using Hölder’s inequality and $\Gamma(z)$ instead of $\Gamma$:
$\|a\|_{\infty}^{2}\,\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\left(\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right)^{2}\right)^{\frac{1}{2}}.$ (6.10)
Focus on the double integral in the second factor. Using Hölder’s inequality to sum and using the bound observed for $\Gamma(z)$ gives the following upper bound for it:
$2\pi\,\int_{D\setminus D(4T)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz.$
By using that $d(\gamma z,z^{\prime})$ equals $d(z,\gamma^{-1}z^{\prime})$ and interchanging the order of the sum and the integral using the linearity of integral, the previous expression equals
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,\gamma^{-1}z^{\prime})}^{2}dz^{\prime}dz,$
which further equals under a change of variables $z^{\prime}\mapsto\gamma z^{\prime}$ the following:
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Due to the presence of $F_{t,t^{\prime},d(z,z^{\prime})}$, for non-zero contribution, we must have that $d(z,z^{\prime})$ is less than $2T$. This means that the previous expression equals
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)\cap B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz,$
which is then bounded by
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
As the $\gamma$-dependency is now eliminated, the sum contributes the number of elements it sums over as a coefficient, which leads to the following equal expression:
$4\pi^{2}\,\frac{e^{8T}}{\operatorname{InjRad}_{X}^{2}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Hence we get the following bound for Expression (6.10), as now we have a bound for the second factor:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
By writing $z^{\prime}$ using polar coordinates around $z$, whence $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, we get the following equal expression:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz,$
which then equals
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
as there was no $\theta$ dependency nor $z$ dependency.
The following lemma bounds a complicated integral in which $F_{t,t^{\prime},\rho}$ appears. The bounds for $F_{t,t^{\prime},\rho}$ are established in Section 7.
Lemma 6.6.
Let $\beta$ be a positive constant that is smaller than $1$. Then:
$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where
$F_{t,t^{\prime},\rho}\leq 32\pi\,\cdot\,\begin{cases}&\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad\text{ if }\rho<t-t^{\prime}\\ &\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},\qquad\text{ if }t-t^{\prime}<\rho<t\\ &\frac{1}{\sqrt{\sinh(\rho)}},\qquad\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
and where
$C_{6}=480\pi.$
The proof uses bounds for $F_{t,t^{\prime},\rho}$. Acquiring good enough bounds for the integral to converge requires very careful analysis which is done in Section 7. There we acquire three different bounds for $F_{t,t^{\prime},\rho}$ working in separate cases. Hence to bound the integral, we split it into three cases and use the bounds for $F_{t,t^{\prime},\rho}$. One of the cases is more technical as both the Beta function and the Gamma function appear in the analysis.
Proof.
We have that
$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{-\frac{\beta}{2}\rho},$
which can be verified by studying the derivatives of the functions. Using this observation gives the following upper bound for the integral we are bounding:
$\frac{1}{\beta}\,\int_{0}^{2T}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}\,F_{t,t^{\prime},\rho}^{2}d\rho.$
By splitting the integral and using the bounds for $F_{t,t^{\prime},\rho}$, we have the following upper bound for the integral expression we set out to bound:
$\frac{32\pi}{\beta}\left(\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{\sqrt{\sinh(\rho)}\,e^{-\frac{\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\ +\ \int_{t-t^{\prime}}^{t}\frac{e^{-\frac{\beta}{2}\rho}}{(1-e^{t-t^{\prime}-\rho})^{\frac{1}{4}}}d\rho\ +\ \int_{t}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.11)
Let us focus on the first term inside the parentheses. It is bounded by
$\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{e^{\frac{1-\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\leq\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\int_{0}^{t-t^{\prime}}\frac{1}{\sqrt{t-t^{\prime}-\rho}}d\rho,$
where the inequality followed from using the upper boundary of the integration domain. The previous then equals after evaluating the integral:
$2\,\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\sqrt{t-t^{\prime}}\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
The inequality followed from the observation that a positive $x$ divided by $\sinh(x)$ is bounded from above by $1$.
Next, let us bound the second term inside the parentheses of Expression (6.11). We begin by writing it in a different form using a change of variables $u=t^{\prime}+\rho-t$:
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{t^{\prime}}e^{-\frac{\beta}{2}u}\,\left(1-\frac{1}{\cosh(u)}\right)^{-\frac{1}{4}}\,du.$
We do another change of variables $z=e^{-u}$ to get
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,\frac{dz}{z},$
where especially
$1-\frac{1}{\cosh(u)}=\frac{(1-z)^{2}}{1+z^{2}}.$
The previous integral expression equals
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}-1}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,dz\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{1}z^{\frac{\beta}{2}-1}\,(1-z)^{-\frac{1}{2}}\,dz,$
where the inequality followed from that $(1+z^{2})^{\frac{1}{4}}$ is bounded by $2$ and the integration domain is enlargened. The previous expression then equals
$2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,B\left(\frac{\beta}{2},\frac{1}{2}\right),$ (6.12)
where $B$ is the Beta function which can be written using Gamma functions (the reference for the Beta and Gamma functions appearing here is [Olv10, Chp.6]):
$B\left(\frac{\beta}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{1}{2}\right)\,\Gamma\left(\frac{\beta}{2}\right)}{\Gamma\left(\frac{1}{2}+\frac{\beta}{2}\right)}.$
We have that $\Gamma\left(\frac{1}{2}\right)$ equals $\sqrt{\pi}$. We also have that
$\Gamma(x)=\frac{1}{x}\Gamma(x+1),$
and that
$\frac{\Gamma(x+1)}{\Gamma\left(x+\frac{1}{2}\right)}\leq(x+1)^{\frac{1}{2}},$
for positive $x$. Hence
$B\left(\frac{\beta}{2},\frac{1}{2}\right)\leq\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}.$
Hence Expression (6.12) is bounded by
$2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
Let us focus then on the third term inside the parentheses of Expression (6.11). The integral can be evaluated thus giving the following equal expression:
$\frac{2}{\beta}\left(e^{-\frac{\beta}{2}t}-e^{-\frac{\beta}{2}(t+t^{\prime})}\right),$
which has the following upper bound
$\frac{2\,e^{-\frac{\beta}{2}(t-t^{\prime})}}{\beta}.$
All three terms inside the parentheses of Expression (6.11) are bounded and they give the final upper bound for the original integral:
$\frac{32\,\pi}{\beta}\,e^{-\frac{\beta}{2}(t-t^{\prime})}\left(2\ +\ 2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\ +\ \frac{2}{\beta}\right)\leq\frac{480\pi}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where we especially noted that $\beta<1$ for the inequality.
7. Weight function
Consider the following integral:
$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}},$
where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. The following lemma shows that the chosen points do not matter as long as the distance between the points is the same.
Lemma 7.1.
Let $z,z^{\prime},w,w^{\prime}\in\mathbb{H}$ such that $d(z,z^{\prime})$ equals $d(w,w^{\prime})$. Then
$\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}$
$=\ \int_{B(w,t)\cap B(w^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}}.$ (7.1)
The proof is based on using the properties of the isometries in the hyperbolic plane.
Proof of Lemma 7.1.
There exists an isometry $f$ such that $f(z)$ equals $w$ and that $f(z^{\prime})$ equals $w^{\prime}$. Using a change of variables $x\mapsto f^{-1}(x)$, we get the following equal expression for the left-hand side of Expression (7.1):
$\int_{f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(f^{-1}(x),z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(f^{-1}(x),z^{\prime}))}}$
$=\ \int_{f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}},$
where the equality followed from the fact that $d(f^{-1}(x),z)=d(x,f(z))$ which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$. As $f$ maps $z$ to $w$ and $z^{\prime}$ to $w^{\prime}$, we have that
$f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)=B(w,t)\cap B(w^{\prime},t^{\prime}).$
Using the previous observation gives us that the previous integral expression equals
$\int_{B(w,t)\cap B(w^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}}.$
The main goal of this section is to give an upper bound for $F_{t,t^{\prime},\rho}$. Here, we additionally assume that $t>t^{\prime}$ for the case $t^{\prime}<t$ is symmetric. $F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data appearing in the proof of Theorem 1.5 in Proposition 6.1. Bounding this weight is done in three lemmas, Lemmas 7.3, 7.4, and 7.5, of which settings are illustrated, respectively, in the following three figures:
Figure 2. The case of $t>t^{\prime}+\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ for it is enveloped by the bigger ball $B(z,t)$. The bound in this case is established using the polar coordinates around $z^{\prime}$.
Figure 3. The case of $\rho<t<t^{\prime}+\rho$, where $a+b=\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is contained in the bigger ball. The integration domain is split into two parts, $D_{1}$ and $D_{2}$ by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to points $w$ and $w^{\prime}$ which are the points where the two balls intersect. The bound in this case is established separately for $D_{1}$ and $D_{2}$. For $D_{1}$, we use the polar coordinates around $z^{\prime}$ and for $D_{2}$ the polar coordinates around $z$.
Figure 4. The case of $t<\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball. The bound in this case is established using the polar coordinates around $m$ which is the intersection of geodesics $z,z^{\prime}]$ and $w,w^{\prime}]$, where $w$ and $w^{\prime}$ are the intersections of the two balls.
From these three lemmas we get the following bound for $F_{t,t^{\prime},\rho}$.
Proposition 7.2.
$F_{t,t^{\prime},\rho}\leq C_{1}\,\cdot\,\begin{cases}\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad&\text{ if }\rho<t-t^{\prime}\\ \frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{\frac{1}{8}},\qquad&\text{ if }t-t^{\prime}<\rho<t\\ \frac{1}{\sqrt{\sinh(\rho)}},\qquad&\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
where
$C_{1}=\max\left(C_{2},C_{3},C_{4}\right)=\max(32,32\pi,99)=32\pi.$
The following lemma gives the bound for $F_{t,t^{\prime},\rho}$ in the first case, see Figure 2.
Lemma 7.3.
Assume that $t$ is larger than $t^{\prime}+\rho$. Then
$F_{t,t^{\prime},\rho}\leq\frac{C_{2}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where
$C_{2}=32.$
The proof of lemma is based on using two different lower bounds for $\cosh(t)-\cosh(R)$, when working in the polar coordinates around $z^{\prime}$. Using these two bounds simultaneously allows us to deal with the problematic small angles, when the boundaries of two balls are close to each other.
Proof of Lemma 7.3.
We write the integral using the polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$. Hence $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$. The situation is illustrated in Figure 2. Hence $F_{t,t^{\prime},\rho}$ is written as
$\int_{0}^{t^{\prime}}\int_{-\pi}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
$=\qquad$ $2\int_{0}^{t^{\prime}}\int_{0}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}},$ (7.2)
where half of the angles are dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ using the polar coordinates $r$ and $\theta$.
We want to find a lower bound for $\cosh(t)-\cosh(R)$. We remark that we have the following inequality since $t>t^{\prime}+\rho>R$:
$\cosh(t)-\cosh(R)\geq\cosh(t)-\cosh(t^{\prime}+\rho).$
This implies the following inequality by using the mean value theorem and the convexity of the hyperbolic sine for the positive input:
$\cosh(t)-\cosh(R)\geq\sinh(t^{\prime})\,\sinh(\rho)\,(t-t^{\prime}-\rho).$ (7.3)
We can also get another inequality. We write $\cosh(t)$ as follows:
$\cosh(t)=\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime}).$
Then, by using the hyperbolic law of cosines, we have that
$\cosh(R)=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$
Since $r<t^{\prime}$ and $\rho<t-t^{\prime}$, we have the following inequality:
$\cosh(t)-\cosh(R(r,\theta))\geq(1-\cos(\theta))\,\sinh(t)\,\sinh(t-t^{\prime}).$ (7.4)
We write Expression (7.2) as
$2\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}d\theta dr,$
which has the following upper bound
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(1-\cos(\theta))^{\frac{1}{4}}}d\theta dr,$
which was acquired by using the inequalities appearing in Expressions (7.3) and (7.4). The following equal expressions are acquired by evaluating the integrals:
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,2^{\frac{5}{4}}\,K\left(\frac{1}{2}\right)dr$
$=\ \frac{2^{2+\frac{5}{4}}\,K(\frac{1}{2})\,\sqrt{\cosh(t^{\prime}-1)}}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where $K$ is the complete elliptic integral of the first kind (see [Olv10, Chp.19]). The previous expression has a further upper bound, where we also bound the elliptic integral by $2$:
$\frac{2^{\frac{17}{4}}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\leq 32\,\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}.$
We proceed to a lemma bounding the second case, see Figure 3.
Lemma 7.4.
Assume that $\rho<t<\rho+t^{\prime}$. Then:
$F_{t,t^{\prime},\rho}\leq C_{3}\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},$
where
$C_{3}=32\pi.$
The idea of the proof is to split the integration domain into two separate domains. The bounding of the other domain resembles the approach of the previous lemma in that we establish two different lower bounds for $\cosh(t)-\cosh(R)$, which allow us to get a bound when used simultaneously. Bounding the other domain turns out to be simple. The following figure, Figure 5, further illustrates these separate domains:
((a)) The integration domain $D_{1}$ consists of those points that are outside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, the polar coordinates around $z^{\prime}$ are used.
((b)) The integration domain $D_{2}$ consists of those points that are inside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, polar coordinates around $z$ are used.
Figure 5. The case is split into two case which are bounded separately.
Proof of Lemma 7.4.
The integration domain of $F_{t,t^{\prime},\rho}$ can be split into a sum of integration over two disjoint domains, $D_{1}$ and $D_{2}$, see Figure 3. We bound these separately. $D_{1}$ is better illustrated in Figure 5(a) and $D_{2}$ in Figure 5(b).
Focus on $D_{1}$ first. We write it using the polar coordinates around $z^{\prime}$. Thus $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$, where $\theta=0$ corresponds to direction $z^{\prime},z](z^{\prime})$. In this way, the integral over $D_{1}$ equals
$\int_{-\Theta}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
= $=\ 2\int_{0}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
where half the angles were dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ in the polar coordinates $r$ and $\theta$. Notice that the maximum angle $\Theta$ is larger than $\pi/2$. We split the integral into two terms as follows:
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
+ $+\ 2\int_{\frac{\pi}{2}}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$ (7.5)
By using the hyperbolic law of cosines we get that
$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})-\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.6)
$\cosh(R)=\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$ (7.7)
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\cos(\theta)\sinh(r)-\cos(\Theta)\sinh(t^{\prime})).$ (7.8)
Hence we have the following inequalities:
$\cosh(t)-\cosh(R)\geq\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))$ (7.9)
and
$\cosh(t)-\cosh(R)\geq\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r)).$ (7.10)
Using the inequality of Expression (7.9), we get that the first term of Expression (7.5) is bounded by
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
By evaluating the integral over $r$, we get the following equal expression:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)-\cos(\Theta)}}\,d\theta.$
This has the following upper bound, when we notice that both $\cos(\theta)$ and $\cos(\Theta)$ are positive. Additionally the ratio with the hyperbolic sine and cosine is bounded by $1$:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)}}\,d\theta.$
The following bound follows from the well known upper bound for the integral:
$\frac{12}{\sqrt{\sinh(\rho)}}.$
Let us focus on the second term of Expression (7.5). We write it as follows by interchanging the integrals using Fubini’s theorem:
$2\,\int_{0}^{t^{\prime}}\int_{\frac{\pi}{2}}^{\Theta}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}\,d\theta\,dr.$
We use the two inequalities appearing in Expressions (7.9) and (7.10) to get the following upper bound for the previous expression:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\cos(\theta)-\cos(\Theta))^{\frac{1}{4}}}\,d\theta\,dr.$
Using the mean value theorem, we get the following upper bound:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\Theta-\theta)^{\frac{1}{4}}}d\theta dr.$
Integrating the inner integral yields
$\frac{4}{3}\,\left(\Theta-\frac{\pi}{2}\right)^{\frac{3}{4}}.$
Hence we have the following bound for the previous integral expression:
$\frac{4\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}dr$
$=\ \frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}},$ (7.11)
where the equality followed from the evaluation of the remaining integral.
Focus momentarily on $\sin(\Theta)$. Expression (7.6) can be modified into the following form using the hyperbolic identities:
$\cos(\Theta)=1-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}.$
The subtracted division term must be larger than $1$ as the cosine is negative due to $\Theta$ being in $(\frac{\pi}{2},\pi)$. We have that
$\sin(\arccos(1-x))=\sqrt{x}\sqrt{2-x}\geq\sqrt{2-x},$
when $x>1$. Hence:
$\sin(\Theta)\geq\sqrt{2-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}}=\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\sinh(t^{\prime})\sinh(\rho)}},$
where the equality followed using hyperbolic identities. The previous implies the following inequality:
$\sin(\Theta)\geq\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\cosh(t^{\prime}+\rho)}}=\sqrt{1-\frac{\cosh(t)}{\cosh(t^{\prime}+\rho)}}.$
This then gives us the following inequality:
$\sin(\Theta)\geq\sqrt{1-\frac{1}{\cosh(t^{\prime}+\rho-t)}}.$ (7.12)
Applying the inequality of Expression (7.12) in Expression (7.11) gives the following upper bound:
$\frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
This has the following further upper bound:
$\frac{16\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The integration over $D_{1}$ has the following upper bound, when both of the terms of Expression (7.5) are added together:
$\frac{1}{\sqrt{\sinh(\rho)}}\left(12+16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\right)$
$\leq\ 16\pi\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
Next, we will focus on bounding the integral over $D_{2}$. We write it using the polar coordinates around $z$. Thus $x=\pi_{1}(\phi_{R}(z,\theta))$, where $\theta=0$ corresponds to direction $z,z^{\prime}](z)$. Then the integral equals:
$\int_{-\Psi}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta$
= $=\ 2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta,$
where the equality followed from using the symmetry. $r=r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates $R$ and $\theta$. $R(\theta)$ denotes the minimum distance that needs to be propagated into direction $\theta$ so that we are in $D_{2}$. Notice also that $\Psi$ is less than $\pi/2$. By using the hyperbolic law of cosines we have that
$\cosh(t^{\prime})=\cosh(\rho)\cosh(t)-\sinh(\rho)\sinh(t)\cos(\Psi),$
$\cosh(r)=\cosh(\rho)\cosh(R)-\sinh(\rho)\sinh(R)\cos(\theta).$
As $\cosh(r)$ grows as $R$ grows, we have the following lower bound for $\cosh(t^{\prime})-\cosh(r)$:
$\sinh(\rho)\sinh(t)(\cos(\theta)-\cos(\Psi)).$
Hence the integral expression has the following upper bound:
$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\frac{1}{\sqrt{\sinh(t)}}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}}dRd\theta.$
This has an upper bound when we integrate from 0 instead of $R(\theta)$; the following expression is acquired after we evaluate the integral:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}d\theta.$
This is then bounded by
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}d\theta,$
which has the following upper bound using the mean value theorem:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
Using the well-known bound for sine, we get the following upper bound:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\frac{2}{\pi}\theta}}d\theta.$
This is actually the Beta integral (see [Olv10, Chp. 5]), so we get the following equal expression:
$\frac{4\sqrt{\pi}\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}}.$
The integration over both $D_{1}$ and $D_{2}$ is done. Uniting the bounds of these two cases gives us the final bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\left(16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\ +\ \frac{4\pi\sqrt{\pi}}{\sqrt{2}}\right)$
$\leq\qquad$ $\frac{32\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The following lemma takes care of the last case, see Figure 4.
Lemma 7.5.
Assume that $t>\rho$. Then:
$F_{t,t^{\prime},\rho}\leq C_{4}\,\frac{1}{\sqrt{\sinh(\rho)}},$ (7.13)
where
$C_{4}=99.$
The proof is again based on splitting the integration domain into two. This time the splitting is done by $w,w^{\prime}]$ which splits the integration into two sides, one closer to $z$ and one closer to $z^{\prime}$. The technique used to bound both sides is the same, based on establising carefully bounds using hyperbolic geometry.
Proof of Lemma 7.5.
The integration domain of $F_{t,t^{\prime},\rho}$ is the intersection of $B(z,t)$ and $B(z^{\prime},t^{\prime})$. Since $t<\rho+t^{\prime}$, the intersection is not a ball. Let $w$ and $w^{\prime}$ be the intersections of the boundaries of the balls, and let $m$ be the intersection of $z,z^{\prime}]$ with $w,w^{\prime}]$. We parametrize the domain using the polar coordinates around $m$ with $\theta=0$ corresponding to direction $z,z^{\prime}](m)$. Hence $x=\pi_{1}(\phi_{r}(m,\theta))$. To easen the notation, we call $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$. We will present $d(z,x)$ and $d(z^{\prime},x)$ using polar coordinates, denoted by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectfully. For illustration, see Figure 4. Hence $F_{t,t^{\prime},\rho}$ is in polar coordinates
$\int_{0}^{2\pi}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}},$
where $r(\theta)$ is the maximum distance we can go from $m$ to direction $\theta$ to still stay in the integration domain. We can split the integration into two parts:
$\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}$
$+\ \int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}.$ (7.14)
The first term integrates over the side closer to $z^{\prime}$ while the second term integrates the side closer to $z$, these sides are separated by $w,w^{\prime}]$.
Let us focus on the integrand of the first term. Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where we note that $\cos(\theta)$ is positive. We also notice that by using the hyperbolic law of cosines we get that
$\cosh(t)=\cosh(a)\cosh(h),$
$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
which implies that
$\cosh(t^{\prime})=\frac{\cosh(b)}{\cosh(a)}\cosh(t).$
Inserting the earlier equality of $\cosh(t)$ into the previous yields the following:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\sinh(a)\left(\sinh(r(\theta))-\sinh(r)\right),$
which then has a lower bound:
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
and then has a lower bound
$\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))$ is bounded from below by
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
This equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(b)\sinh(a)\sinh(r(\theta))+\cosh(a)\sinh(b)\sinh(r)\right),$
and is then bounded from below by
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(b)\sinh(a)+\cosh(a)\sinh(b)\right).$
As $a+b=\rho$, the previous equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
Now, let us focus on the integrand of the second term of Expression (7.14). Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))-\sinh(b)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where $\cos(\theta)$ is negative. We also notice that, using the hyperbolic law of cosines,
$\cosh(t)=\cosh(a)\cosh(h),$
$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
which implies that
$\cosh(t)=\frac{\cosh(a)}{\cosh(b)}\cosh(t^{\prime}).$
The previous gives the following equality after inserting an earlier equality of $\cosh(t^{\prime})$ in:
$\cosh(t)=\cosh(a)\cosh(r(\theta))-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)-\cos(\theta)\sinh(b)\left(\sinh(r(\theta))+\sinh(r)\right).$
This has a lower bound
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))-\sinh(a)\sinh(r)\right),$
and then has a lower bound
$(-1)\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R(r,\theta)))(\cosh(t^{\prime})-\cosh(P(r,\theta)))$ is bounded from below by
$(-1)\,\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)\,\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
This equals
$(-1)\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(a)\sinh(b)\sinh(r(\theta))+\cosh(b)\sinh(a)\sinh(r)\right),$
and is then bounded from below by
$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(a)\sinh(b)+\cosh(b)\sinh(a)\right).$
As $a+b=\rho$, the previous equals
$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
These observations for the integrands of the terms of Expression (7.14) imply that we get the following upper bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta$ (7.15)
Focus on the inner integral. Assume first that $r(\theta)<2$. The integral is bounded by the following by using the mean value theorem:
$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}.$
This is bounded by $2\sqrt{2}$. Then, assume that $r(\theta)>2$. We split the integration as follows:
$\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$
$a$ is specified later. We have the following upper bound for the previous:
$\frac{1}{\sqrt{\sinh(a)}}\,\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr.$
The first term was bounded using the boundary of the integration domain and the second term was bounded using the mean value theorem. By evaluating the integrals, we acquire the following equal expression:
$\frac{2\,\sqrt{\cosh(r(\theta))-\cosh(a)}}{\sqrt{\sinh(a)}}\ +\ 2\,\sqrt{r(\theta)-a}\leq\frac{2\,\sqrt{r(\theta)-a}\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta))}}\ +\ 2\,\sqrt{r(\theta)-a}.$
The inequality followed from using the mean value theorem. We choose $a$ to equal $r(\theta)-1$. Then the previous equals
$\frac{2\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta)-1)}}\ +\ 2\leq\frac{2\,\sqrt{\sinh(2)}}{\sqrt{\sinh(1)}}\ +\ 2,$
where for the inequality we used that $r(\theta)>2$. The previous is bounded by $9$.
The previous observations mean that Expression (7.15) has the following upper bound:
$\frac{9}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta.$
The integral is bounded from above by $11$, so we can conclude that
$F_{t,t\,\rho}\leq\frac{99}{\sqrt{\sinh(\rho)}}.$
8. Proof of main theorems
In this section we give proofs to the main theorems of this paper, Theorems 1.4, 1.5, and 1.7. We first prove the quantitative main theorem, Theorem 1.5. Most of the paper is dedicated to this end. Proving it is highly technical and thus for the clarity of the proof we have presented auxiliary results in separate sections, Sections 4, 5, 6, and 7.
After proving Theorem 1.5, we prove the limit form main theorem, Theorem 1.4. It follows straightforwardly from Theorems 1.5 and 8.1.
Proving Theorem 1.7 relies on using Theorems 1.5 and 8.2. The proof is very similar to that of Theorem 1.4 but due to the probabilistic setting imposing lighter conditions on the sequence, we need Theorem 8.2 to record some key results in random surface theory.
We proceed to prove Theorem 1.5. We restate the theorem for the convenience of the reader.
Theorem.
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
$0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
and fix $\tau\in\mathbb{R}$. Then
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The coefficient appearing is
$C_{8}=50000\pi.$
The proof presented here is clear as the technical details are dealt in different sections, see Sections 6, 5, 6, and 7. The core idea of the proof is to split the problem into a geometric side and a spectral side using a well-chosen propagator. The geometric side and the spectral side can then be dealt separately.
Proof of Theorem 1.5.
Define functions
$h_{t}(x)=x^{-1}\sin(tx),\qquad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x),$
and operators
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right).$
Then due to way the operators are defined, we have that
$P_{t}\psi_{j}=h_{t}(\rho_{j})\psi_{j},\qquad P_{t,\tau}\psi_{j}=h_{t,\tau}(\rho_{j})\psi_{j}.$
Using the previous, we can write that
$|\langle\psi_{j},a\psi_{k}\rangle|=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|^{-1}\left|\left\langle\psi_{j},\,\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\,\psi_{k}\right\rangle\right|.$
The left factor containing spectral data is bounded using Proposition 4.1,which can be used since $I$ is compact in $(0,\infty)$, $T\delta=\pi/2$, and $\delta$ is smaller than $(2/9)\cdot\min(I)$:
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})\,dt\right|^{-1}\leq 8\pi\rho_{j}\,\rho_{k}\leq 8\pi\max(I)^{2}.$
Thus,
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}|\langle\psi_{j},a\psi_{k}\rangle|^{2}$ $\leq 64\pi^{2}\,\max(I)^{4}\,\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\psi_{k}\right\rangle\right|^{2}$
$\leq 64\pi^{2}\,\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2},$
where the Hilbert-Schmidt norm bound is acquired when we consider all the basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
Next we want to apply Proposition 6.1 to bound the Hilbert-Schmidt norm. This requires understanding the integral kernels of $P_{t}$ and $P_{t,\tau}$ for which a convenient form is given in Proposition 5.1. We apply Proposition 6.1 to acquire the following upper bound for the previous expression:
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we opted to use $|X\setminus X(4T)|$ to denote the volume of the points of $X$ with injectivity radius smaller that $4T$. Also,
$C_{7}=\frac{7200}{\pi}.$
Hence we can conclude that
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$C_{7}\cdot 64\pi^{2}\leq 50000\pi=C_{8}.$
We now turn to Theorem 1.4, the limit form main theorem. We also restate the theorem for convenience.
Theorem.
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Assume that the sequence has the following properties:
  • (BSC)
    Benjamini–Schramm convergence: for every $R>0$,
    $\lim_{n\to\infty}\frac{\big|\{x\in X_{n}:\operatorname{InjRad}_{X_{n}}(x)<R\}\big|}{|X_{n}|}=0,$
  • (EXP)
    Expander property: the sequence admits a uniform lower bound on the spectral gap,
  • (UND)
    Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
  • (i)
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon.$
Properties (ii) and (iii) remain to be proven; Property (i) was established by Le Masson and Sahlsten (see Theorem 1.1 [LS17]). We remark that the proof presented here for properties (ii) and (iii) could be adapted with slight additions to also include property (i), thus giving a new proof for the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law which is used to asymptotically relate the volume of the surface with the number of eigenvalues in an energy window, when we proceed in the sequence of surfaces. Before the proof, we state the large-scale Weyl law (see [LS17, Sec.9]).
Theorem 8.1 (Large-scale Weyl law).
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying property (BSC) stated in Theorem 1.4. $X_{n}$ has Laplace-Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
Then
$\frac{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}{|X_{n}|}=\frac{\#\{j\ :\ \lambda_{j}^{(n)}\in J\}}{|X_{n}|}\xrightarrow{n\to\infty}C,$
where $C$ is a constant depending on the sequence of surfaces and the interval $I$.
We have included using both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using $\lambda_{j}^{(n)}$ approach. However, the theorem works just as well for $\rho_{j}^{(n)}$ as was seen in the statement of the previous theorem.
We proceed to prove Theorem 1.4.
Proof of Theorem 1.4.
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)>0$ such that, for all $\tau\in\mathbb{R}$,
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\sum_{j:\rho_{j}^{(n)}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle\right|^{2}<\epsilon,$ (8.1)
where
$N(X_{n},I)=\#\{j\ :\ \rho_{j}^{(n)}\in I\}.$
Let us define a following function:
$\bar{a}_{n}=a_{n}-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx.$
$\bar{a}_{n}$ is a mean 0 normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have that
$\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle=\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle.$
Hence we have the following equal expression for the left-hand side of Expression (8.1):
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\sum_{j:\rho_{j}^{(n)}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle\right|^{2}$
We can apply Theorem 1.5 to the double sum given that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius being smaller than $4T$, $T\,\delta=\pi/2$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{8}=50000\pi.$
Then we apply the properties of the sequence to tidy the previous expression. First we note that
$\|\bar{a}_{n}\|_{2}\leq\sqrt{|X_{n}|}\,\|\bar{a}_{n}\|_{\infty}.$
And then, since $(a_{n})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{n}\|_{\infty}\leq 2\,a_{max}.$
Property (EXP) means that there is a universal lower bound for the spectral gaps which then implies that $\beta(\lambda_{1}^{(n)})$ has a universal lower bound. Denote this by $\beta_{min}$. Finally, Property (UND) means that there is a universal lower bound for the injectivity radii. Denote this by $\operatorname{InjRad}_{min}$. Using these observations gives the following upper bound for Expression (8.2):
$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{8\pi^{2}\,|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{8\pi^{2}\,e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
where the equality followed from splitting the limit supremum into two terms.
Consider the second limit supremum inside the parentheses of the previous expression. Due to the large-scale Weyl law, Theorem 8.1, we have that $N(X_{n},I)$ grows asymptotically like the volume of surfaces $|X_{n}|$. This means that asymptotically the ratio inside the limit supremum closes to 0 as we proceed in the sequence since Property (BSC) forces the ratio of the volume of the points with small injectivity radius and the volume of the surface to go to zero. Hence the previous expression equals
$\frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right).$
Using the large-scale Weyl law again to the remaining limit supremum, we see that the ratio approaches some constant $C_{9}$. Hence the previous expression equals
$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}$
$=\ \frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
where for the equality we used that $T\,\delta=\pi/2$. If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{8}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
In the end we prove the probabilistic theorem, Theorem 1.7. We restate the theorem for convenience.
Theorem.
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
Properties (ii) and (iii) remain to be proven; Property (i) was already proven by Le Masson and Sahlsten [LS24]. The proof is very close to that of Therem 1.4. Now we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced. This is done by the following theorem acquired from the theory of the Weil-Petersson random surfaces.
Theorem 8.2.
Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil-Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with associated values
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}.$
Then with probability 1o(1)1-o(1) $X$ has the following properties:
  • (a)
    $\operatorname{InjRad}_{X}\geq g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}},$
  • (b)
    $\frac{|\{x\in X\ :\ \operatorname{InjRad}_{X}(x)<\frac{1}{6}\,\log(g)\}|}{|X|}=\mathcal{O}(g^{-\frac{1}{4}}),$
  • (c)
    $\lambda_{1}\geq\frac{1}{4}-\eta,$
where $\eta\in(0,\frac{1}{4})$. Additionally,
  • (d)
    $\{j\ :\ \rho_{j}\in I\}=\#\{j\ :\ \lambda_{j}\in J\}=|X|\cdot\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right),$
where $I\subset(0,\infty)$ is a compact interval and
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
The theorem is acquired from two articles: [Mon22, HMT25]. Properties (a), (b), and (d) together with the rates is acquired from [Mon22], and Property (c) together with rate is acquired from [HMT25]. When comparing the proof of Theorem 1.7 with the proof of Theorem 1.4, Property (a) replaces Property (UND), Property (b) replaces Property (BSC), and Property (c) replaces Property (EXP). Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten working in this probabilistic setting.
Proof of Theorem 1.7.
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have that
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
where
$N(X_{g},I)=\#\{j\ :\ \rho_{j}^{(g)}\in I\}.$
Let us define a following function:
$\bar{a}_{g}=a_{g}-\frac{1}{|X_{g}|}\int_{X_{g}}a_{g}\,dx.$
$\bar{a}_{g}$ is a mean 0 normalization of $a_{g}$. Due to orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have that
$\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle=\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle.$
Hence the left-hand side of Expression (8.3) equals
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
We can apply Theorem 1.5 to the double sum givem that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\,\delta=\pi/2$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{8}=50000\pi.$
First notice that
$\|a_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|a_{g}\|_{\infty}.$
Since the sequence $(a_{g})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{g}\|\leq 2\,a_{max}.$
Hence the previous limit supremum expression is bounded from above by
$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
Using Theorem 8.2, we have that
$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
$\leq\ \frac{1}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(\frac{3}{16}-\eta\right)^{3}}\,\left(|X_{g}|+8\pi^{2}\,|X_{g}|\,\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
with propability $1-\mathcal{O}(g^{-c(\eta)})$, where $\eta\in(0,3/16)$, $c(\eta)$ is a positive constant depending on $\eta$, and where
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
The previous expression has a tidier upper bound:
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+8\pi^{2}\,e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
This means that with high propability Expression (8.4) equals
$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta$
where equality followed from using that $T\,\delta=\pi/2$. $C_{10}$ is a constant upper bound coming from
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}.$
If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
Acknowledgements.
The author would like to thank Prof. Mostafa Sabri, Prof. Tuomas Sahlsten, and Prof. Farrell Brumley for discussions that led to this article. The author also thanks Prof. Tuomas Sahlsten, Prof. Henrik Ueberschär, Dr. Félix Lequen, Dr. Søren Mikkelsen, and Dr. Joe Thomas for their guidance while writing this article.
Our use of the exponential is inspired by the upcoming article [HLM+] of the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten and Henrik Ueberschär on quantum ergodicity for eigenfunctions for Schrödinger operators on hyperbolic surfaces. We thank the other authors of that paper.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University.